OpenStudy (anonymous):

(6-y/y-2)-(8y-7/2-y)

OpenStudy (anonymous):

Isn't this the same question you posted before?

OpenStudy (anonymous):

\[\ \frac{6 - y}{y-2} \ - \ \frac{8y - 7}{2 - y} \]

OpenStudy (anonymous):

You have to understand that before you can add any two fractions their denominators have to be the same. So in this case what we can do is to rewrite the denominator of the right hand expression like this: -(y - 2)...Multiplying the a negative reverses the +/- signs in the parentheses. Now you must also realize that our operation will be addition because of this new negative times the one we already have there. So we have: \[\ \frac{6-y}{y-2} \ + \ \frac{8y-7}{y-2} \]

OpenStudy (anonymous):

Now all that is left is to add the like terms: \[\ \frac{8y - y+6-7}{y-2} \ = \ \frac{7y-1}{y-2} \]